OpenStudy (princessaurora):

*Will fan and medal* Model each problem as an equation, and then match it to its solution. 1) When the sum of the squares of two positive integers is 185. If one integer is 3 less than the other, find the larger integer. 2) The numerator of a fraction is 1 more than twice its denominator. If 4 is added to both the numerator and the denominator, the fraction reduces to 5/3. Find the denominator. 3) The difference of a positive integer and its inverse is 15/4. Find the integer. *Solutions* A) 11 B) 4 C) 5

OpenStudy (princessaurora):

@netlopes1 help me por favor

OpenStudy (bibby):

3) The difference of a positive integer and its inverse is 15/4. Find the integer. \(a-\dfrac{1}{a}=\dfrac{15}{4}\iff \dfrac{a}{1}-\dfrac{1}{a}=\dfrac{15}{4}\iff \dfrac{a^2}{a}-\dfrac{1}{a}=\dfrac{15}{4}\\\implies \dfrac{a^2-1}{a}= \dfrac{15}{4}\) from there you have 2 things you can solve for \(a^2-1=15\) or \(a=4\) one of them is obvious

OpenStudy (princessaurora):

The first one is a= 4,-4?

OpenStudy (bibby):

c is 4, just 4

OpenStudy (princessaurora):

so 3) is B)?

OpenStudy (bibby):

1) When the sum of the squares of two positive integers is 185. If one integer is 3 less than the other, find the larger integer. \(a^2+b^2=185 \\a=b-3 \\(b-3)^2+b^2=185 \\(b-3)(b-3)=b^2-6b+9 \\b^2-6b+9+b^2=185 \\2b^2-6b+9=185 \\2b^2-6b\cancel{+9-9}=185-9 \\2b^2-6b=176 \\2(b^2-3b)=176 \\b^2-3b-88=0 \\(b+8)(b-11)=0 \\b=-8,11\)

OpenStudy (bibby):

since we want only positive integers we can discard -8

OpenStudy (princessaurora):

ok I see

OpenStudy (princessaurora):

Thanks so much I appreciate it a lot, I got it now! :)

OpenStudy (bibby):

oh there's another one

OpenStudy (bibby):

I have no idea how to do it lol